OpenStudy (logan13):

What is the value of the expression |x| + |y + z| when x = –7, y = 8, and z = –11? What is the value of the expression |x| + |y + z| when x = –7, y = 8, and z = –11? What is the value of the expression |x| + |y + z| when x = –7, y = 8, and z = –11? What is the value of the expression |x| + |y + z| when x = –7, y = 8, and z = –11?

OpenStudy (anonymous):

7 + 8 + 11 = 26

OpenStudy (welshfella):

the vertical lines mean the absolute value of the expression inside them So for example x = -6 then absolute value of x will be 6 ( always positive)

OpenStudy (anonymous):

Oh...so mine is wrong????

OpenStudy (welshfella):

@ilovecandy101 that is wrong yes

OpenStudy (anonymous):

Oh......OK I am not good at math at all.......

OpenStudy (welshfella):

|y + z| = | 8 - 11| = | -3| = 3

OpenStudy (anonymous):

dont say that you just need to study @ilovecandy101

OpenStudy (welshfella):

right

OpenStudy (anonymous):

@TitansGo!! I guess you're right, that is why I am so happy for Open Study! People answer your questions and show you what to do. I am going to try my best to get better at math!!! Keep Calm And Love Math, well at least try!

OpenStudy (anonymous):

thanks and i hope you do well

OpenStudy (anonymous):

:D